OpenStudy (anonymous):

If √6 is the geometric mean between 6 and another number, then the number is 1 2 6

OpenStudy (amoodarya):

\[\frac{a}{\sqrt{6}}=\frac{\sqrt{6}}{b}\\so\\ab=6\\\] one of a,b =6 so other is ...?

OpenStudy (usukidoll):

ab = 6 as in a x b = 6? @amoodarya

OpenStudy (amoodarya):

yes a*b=6 and for example a=6 so ...

OpenStudy (usukidoll):

ah so find an a and b such that it equals to 6 so for your example a = 6 6b=6 solving for b gives me 1 so (6)(1) = 6 6=6

OpenStudy (amoodarya):

\[1,\sqrt{6},6\\or \\6,\sqrt{6},1\]

OpenStudy (usukidoll):

so there's more than one answer to this. \[(\sqrt{6})(\sqrt{6} )=6\] 6=6

OpenStudy (anonymous):

the geometric mean of \(a,b\) is \(\sqrt{ab}\) so here we have \(\sqrt6=\sqrt{6x}\implies x=1\)

OpenStudy (anonymous):

a,b>=0 or a,b<=o